# Dynamics of Two Species with Density-Dependent Interactions in a Mutualistic Context

### v1

@@ 1 Introduction
In natural communities, interactions between individuals from different species are widespread, ranging from antagonistic relationships, such as competition or parasitism, to cooperative ones, such as mutualism (Bronstein,, 2015; Janzen,, 1985). Mutualistic interactions include processes such as seed dispersal (Janzen,, 1985), pollination (Klein et al.,, 2007), resource exchange (van der Heijden et al.,, 2008), and protection (Bronstein,, 2015), where individuals from both species may benefit from the relationship. For example, mycorrhizal fungi provide plants with essential nutrients, while plants provide carbohydrates to the interacting fungi (Bonfante and Genre,, 2010). Similarly, in pollination, plants produce nectar feeding some insects, which in turn transport pollen and facilitate plant reproduction (Sargent and Ackerly,, 2008).
@@
Mutualistic interactions can be either facultative or obligatory. If facultative, the dynamics of each species can benefit from each other, but each species can be maintained in the absence of the other (e.g. generalist pollination, where plants and pollinators can often find alternative partners (Sargent and Ackerly,, 2008)). Otherwise, if mutualistic interactions are obligatory, the persistence of each species entirely depends on the mutualistic relationship and hence on the presence of the interacting species (e.g. the fig-wasp relationship, Bronstein, 2001).
@@
The nature and strength of the interactions can also depend on population densities. Variations in species densities can shift the cost-benefit dynamics of the relationship, consequently shifting interactions along a continuum from mutualism to parasitism. For example, in ant-plant mutualism, plants provide ants with food in exchange for protection. However, at low ant densities, plants may not receive sufficient protection, while at high densities, the cost of supporting the ants may outweigh the protective benefits (Bronstein,, 1994; Holland et al.,, 2002; Janzen,, 1985). Similarly, in pollination systems, the investment of plants in nectar can become inefficient if pollinator densities are too low to ensure effective pollination, or too high, leading to increased competition among pollinators and diminishing returns for the plants (Kawai and Tokeshi,, 2007; Morris et al.,, 2010; Holland et al.,, 2002).
@@
Mathematical models of ecological interactions have mostly focused on antagonistic relationships such as predator-prey and competitive dynamics. While some models addressed mutualism, they often assumed strictly positive interactions, leading to unrealistic outcomes like unbounded population growth or infinite densities (Brauer and Castillo-Chavez,, 2012; Hale and Valdovinos,, 2021; May,, 1972). Indeed, because of the complexity and context-dependence of mutualistic interactions, modeling them as a direct extension of Lotka–Volterra systems overlooks key mechanisms, such as density-dependent saturation, that are essential to preventing unbounded growth. Moreover, the factors shaping the persistence of mutualistic interactions and the transitions between parasitic and mutualistic interactions driven by populations densities variations (i.e., on ecological time scales) are much less documented than in predator–prey systems (Holland et al.,, 2002; Revilla and Encinas-Viso,, 2015). Density dependence and saturation effects to prevent unbounded growth or systematic extinction were considered in Brauer and Castillo-Chavez, (2012), bounded growth terms were added in Holland et al., (2002) to model that mutualistic benefits naturally diminish as population sizes increase, and finally Ruiz-Herrera, (2025) analyzed mutualistic systems with a density-dependent framework, emphasizing the role of benefit–cost ratios in shaping long-term interaction outcomes. Asymmetries in obligate vs. facultative mutualistic interactions among species have also been shown to have a significant influence on system stability and resilience (Bronstein,, 1994; Holland and DeAngelis,, 2010; Neuhauser and Fargione,, 2004). Most of these previous models studying the dynamics of mutualistic interactions rely on ordinary differential equations to describe the conditions for coexistence and stability (see Hale and Valdovinos, (2021) for a review). Other approaches such as network analysis have also been used to explore interspecific relationships within ecological systems (Bascompte et al.,, 2003) and partial differential equations to model the spatial dynamics of mutualistic interactions (Aliyu and Mohd,, 2021).
@@
Examples of mutualistic interactions with density-dependent effect have been documented in natural populations (Bronstein,, 2015). Yet, no general theory has been developed to study the coexistence of mutualistic species while accounting for this important process. As pointed out by Hale and Valdovinos, (2021), it is striking that all mutualistic interactions models studied so far show similar outcomes and behaviors, suggesting that some general properties can be found, independently of the peculiarities of each model. For example, there must be some general conditions for the stable coexistence of the two species when the interaction is density-dependent. Previous models of mutualistic interactions often focus on equilibrium points, typically identifying stable nodes or saddle points as the primary outcomes of these systems with particular modeling choices of the functional forms relating the rate of interactions, the birth and death rates, and the populations densities. While these models provide valuable insights, they rarely consider the possibility of more complex dynamics, such as limit cycles or oscillatory behaviors, therefore preventing general predictions.
By proposing and analyzing a general mathematical model of two-species mutualism, we aim at tackling these limitations, focusing in particular on two main questions:
  1. 1.
    What types of dynamical behaviors, including equilibrium configurations, can be derived from models of mutualistic interactions, and which types are not covered by these models?
  2. 2.
    To what extent does incorporating transitions between mutualism and parasitism allow us to observe previously neglected dynamics?
@@
To address these questions, Section 2 first provides the general definitions of density-dependent effects acting on the dynamics of species involved in ecological interactions, which will be used throughout the paper. We thus introduce a system of differential equations for two interacting populations, incorporating density-dependent interactions, without focusing on specific types of interactions within a particular ecological context and on limited range of parameters. In Section 3, we formulate general and robust assumptions derived from the analysis of deterministic mutualistic models (Theorem 3.1). We then present examples of mutualistic interaction models that, despite aiming to represent different aspects of mutualism or emphasizing certain features over others, fall within the framework of Theorem 3.1. In Section 4, we use the extended definition of mutualistic interactions and identify sufficient conditions for the emergence of oscillatory dynamics. These conditions, often overlooked in previous studies, provide new insights into the potential for stable coexistence or extinction. After discussing our results in Section 5, we present the corresponding proofs in Section 6.
@@ 2.1 General Framework
where $x$ and $y$ represent the densities of each species, $f$ and $g$ describe the density-dependent growth rates of each species. Because we model population sizes, the analysis is confined to the non-negative quadrant of $\mathbb{R}^{2}$. This form is commonly used to describe two-species dynamics without accounting for immigration, as in the models of Hale and Valdovinos, (2021), May, (1972), and Brauer and Castillo-Chavez, (2012).
@@ 2.2 Key Assumptions and Definition
To analyze and interpret the behavior of the system, we follow an approach similar to that used in Brauer and Castillo-Chavez, (2012), relying on the partial derivatives of the growth functions. These derivatives capture how the growth rate of each species responds to changes in the density of both species. For this reason, starting from Section 3.2, we will assume that $f$ and $g$ are $\operatorname{\mathcal{C}^{1}}$ functions i.e. continuous and continuously differentiable. We focus on specific regions of the phase plane defined by the population densities $x$ and $y$. In each region, the sign of the partial derivatives of the interaction terms determines the local effect of one species on the growth rate of the other. As a result, an interaction may be positive in one part of the phase portrait and negative in another, depending on the local population densities. Different regions of the phase portrait may correspond to different interaction regimes, described below:
@@
  • Intraspecific Cooperation: $\frac{\partial f}{\partial x}>0$ and $\frac{\partial g}{\partial y}>0$ describe positive feedback within each species, representing cooperative interactions, where individuals of the same population support each other’s growth.
@@
In contrast to many pre-existing models where these partial derivatives always remain within the same strict sign or even remain constant regardless of species densities (Hale and Valdovinos,, 2021), we allow the sign of these partial derivatives to change. Such assumption allows for transitions from mutualistic (positive value) to parasitic interactions (negative value) between species, depending on their respective densities. We thus give an extended definition of mutualism as follows:
@@
These definitions are illustrated in Figure 1, which presents two different situations, both consistent with mutualism as defined in Definition 2.1. In the first case, a strictly mutualistic region appears, although the condition does not hold for all densities $x$ and $y$. In the second case, the system remains mutualistic overall, but no strictly mutualistic region exists; that is, both populations benefit from the interaction, though not simultaneously.
@@
Figure 1: Illustration of density-dependent effects of mutualism (Definition 2.1) through the signs of cross partial derivatives. In both situations, the interaction is classified as mutualistic according to Definition 2.1, even though the local effects of one species on the other vary. Figure 1(a): $\frac{\partial f}{\partial y}>0$ everywhere, meaning that species $y$ always has a positive effect on the growth of species $x$. The hatched region indicates where $\frac{\partial g}{\partial x}>0$, so that species $x$ also benefits species $y$ locally. This region represents a strictly mutualistic zone. Outside this zone, the effect of species $x$ on $y$ is negative, yet the system still qualifies as mutualistic under Definition 2.1. Figure 1(b): The cross partial derivatives $\frac{\partial f}{\partial y}$ and $\frac{\partial g}{\partial x}$ vary and are positive in distinct, non-overlapping regions: the hatched (resp. dashed) region corresponds to $\frac{\partial g}{\partial x}>0$ (resp. $\frac{\partial f}{\partial y}>0$). Outside these regions, the interaction is locally parasitic.

### v2

@@ 1 Introduction
In natural communities, ecological interactions between individuals from different species are widespread and profoundly shape the dynamics of populations. These interactions can be either antagonistic relationships, such as competition or parasitism, or cooperative ones, such as mutualism (Bronstein, 2015; Janzen, 1985), therefore leading to different demographic outcomes in the populations involved. Mutualistic interactions, where individuals from interacting species may benefit from the relationship, include processes such as seed dispersal (Janzen, 1985), pollination (Klein et al., 2007), resource exchange (van der Heijden et al., 2008), and protection (Bronstein, 2015). For example, mycorrhizal fungi provide plants with essential nutrients, while plants provide carbohydrates to the interacting fungi (Bonfante and Genre, 2010). Similarly, in pollination, plants produce nectar feeding some insects, which in turn transport pollen and facilitate plant reproduction (Sargent and Ackerly, 2008).
@@
Mutualistic relationships can be either facultative or obligatory, hence triggering different consequences on the co-existence and dynamics of interacting populations. If mutualistic interactions are obligatory, the persistence of each species entirely depends on the mutualistic relationship and hence on the presence of the interacting species (e.g. the fig-wasp specialized relationship (Bronstein, 2001)). Otherwise, in facultative interactions, the dynamics of each species can benefit from each other, but each species can be maintained in the absence of the other (e.g. generalist pollination, where plants and pollinators can often find alternative partners (Sargent and Ackerly, 2008)). However, the benefits brought to the individuals from the interacting species are not necessarily symmetrical nor constant.
@@
While the mutualistic interactions are classically assumed to bring only benefits to the individuals involved, they can also bring costs, especially in certain conditions of respective densities. Density-dependent benefits have indeed been documented in several communities. Respective densities of interacting species can shift the cost-benefit dynamics of the relationship, consequently shifting interactions along a continuum from mutualism to parasitism. For example, in ant-plant mutualism, plants provide ants with food in exchange for protection. However, at low ant densities, plants may not receive sufficient protection, while at high densities, the cost of supporting the ants may outweigh the protective benefits (Bronstein, 1994; Holland et al., 2002; Janzen, 1985). Similarly, in pollination systems, the investment of plants in nectar can become inefficient if pollinator densities are too low to ensure effective pollination, or too high, leading to increased competition among pollinators and diminishing returns for the plants (Kawai and Tokeshi, 2007; Morris et al., 2010; Holland et al., 2002). There is thus a wide diversity of mutualistic interactions in the wild, and the dynamics of the different species involved can differ depending on the ecological context. Their consequences on population dynamics have been previously described mostly by mathematical models focusing on specific ecological situations, therefore limiting general predictions on the effects of mutualistic interactions.
@@
Many classical models addressing mutualism assumed strictly positive interactions, leading to unrealistic outcomes like unbounded population growth or infinite densities (Brauer and Castillo-Chavez, 2012; Hale and Valdovinos, 2021; May, 1972). Indeed, because of the complexity and context-dependence of mutualistic interactions, modeling them as a direct extension of Lotka–Volterra systems overlooks key mechanisms, such as density-dependent saturation, that prevent unbounded growth in actual populations. Moreover, the factors shaping the persistence of mutualistic interactions and the transitions between parasitic and mutualistic interactions driven by populations densities variations (i.e. on ecological time scales) are poorly documented in mutualistic interactions, while they are well-known in predator–prey systems modeled using Lokta-Volterra equations (Holland et al., 2002; Revilla and Encinas-Viso, 2015). Brauer and Castillo-Chavez (2012) and Holland et al. (2002) considered mutualistic models with saturation effects preventing either unbounded growth or systematic extinction. More recently Ruiz-Herrera (2025) analyzed mutualistic systems with a density-dependent framework, emphasizing the role of benefit–cost ratios in shaping long-term interaction outcomes. Asymmetries in obligate vs. facultative mutualistic interactions among species have also been shown to significantly affect system stability and resilience in a series of models (Bronstein, 1994; Holland and DeAngelis, 2010; Neuhauser and Fargione, 2004). Most of these previous models studying the dynamics of mutualistic interactions relied on ordinary differential equations to describe the conditions for coexistence and stability (see Hale and Valdovinos (2021) for a review). Other approaches such as network analysis have also been used to explore interspecific relationships within ecological systems (Bascompte et al., 2003) and partial differential equations to model the spatial dynamics of mutualistic interactions (Aliyu and Mohd, 2021).
@@
Despite this diversity of modeling approaches, a general theoretical framework investigating the coexistence and dynamics of mutualistic species while accounting for any type of density-dependent effects is still lacking. At the same time, as emphasized by Hale and Valdovinos (2021), it is striking that all mutualistic interaction models studied so far show similar outcomes and behaviors, suggesting that some general properties can be found, independently of the peculiarities of each model. For example, there must be some general conditions for the stable coexistence of the two species when the interaction is density-dependent. Previous models of mutualistic interactions often focus on equilibrium points, typically identifying stable nodes or saddle points as the primary outcomes of these systems with particular modeling choices of the functional forms relating the rate of interactions, the birth and death rates, and the populations densities. While these models provide valuable insights, they rarely consider the possibility of more complex dynamics, such as limit cycles or oscillatory behaviors, therefore preventing general predictions.
By building and analyzing a mathematical model of two-species mutualism with general equations allowing to cover any kind of density-dependent processes, we aim to overcome these limitations and to clarify the range of dynamical behaviors that can emerge in mutualistic systems, including equilibrium structures. To this end, we adopt a framework that is general in two complementary senses. First, the model proposed here is mathematically general: instead of prescribing specific functional responses or mechanistic assumptions, we rely on qualitative and structural constraints on the growth functions and the signs of their partial derivatives. This enables the model to encompass a wide class of two-species interaction systems, including but not limited to most previously published mutualistic models (see Hale and Valdovinos (2021) and Tables in Section 2.1). Second, the model described here is biologically general: because interactions are represented through density-dependent signs of the cross-partial derivatives, the framework is able to incorporate a broad spectrum of ecological processes such as mutualistic benefits, saturation, costs, and shifts to parasitic effects at high densities. Thus, it provides a unified tool for ecological relationships that do not necessarily remain strictly mutualistic across all possible population densities.
@@
This paper is structured as follows: Section 2 first provides the general definitions of density-dependent effects acting on the dynamics of species involved in ecological interactions, which will be used throughout the paper. We introduce a system of differential equations for two interacting populations, incorporating density-dependent interactions, without focusing on specific types of interactions within a particular ecological context and on limited range of parameters. In Section 3, we formulate general and robust assumptions derived from the analysis of deterministic mutualistic models (Theorem 3.1). We then present examples of mutualistic interaction models that, despite aiming to represent different aspects of mutualism or emphasizing certain features over others, fall within the framework of Theorem 3.1. In Section 4, we use the extended definition of mutualistic interactions and identify sufficient conditions for the emergence of oscillatory dynamics. These conditions, often overlooked in previous studies, provide new insights into the potential for stable coexistence or extinction. After discussing our results in Section 5, we present the corresponding proofs in Section 6.
@@ 2.1 General Framework
where $x$ and $y$ represent the densities of each species, $f$ and $g$ describe the density-dependent growth rates of each species. Because we model population sizes, the analysis is confined to the non-negative quadrant of $\mathbb{R}^{2}$. This form is commonly used to describe two-species dynamics without accounting for immigration (see Thompson et al. (2006)), as in the models studied in May (1972), Brauer and Castillo-Chavez (2012), and reviewed in Hale and Valdovinos (2021). The sections 2.1.1 to 2.1.3 present examples of mutualism models studied in previous works, which are of the form (1). For the sake of readability, we have divided them into three groups according to their underlying ecological mechanisms. They are presented with their corresponding $f$ and $g$ functions from Eq. (1).
2.1.1 Linear–Benefit Mutualism Models
Table 1 compiles mutualism models where benefit increases linearly with partner density and there is no saturation effect.
Model Per-capita growth functions Interpretation Reference
Linear benefit with logistic limitation $\begin{aligned} f(x,y)&=r_{1}\left(\dfrac{K_{1}+\alpha_{12}y-x}{K_{1}}\right)\\ g(x,y)&=r_{2}\left(\dfrac{K_{2}+\alpha_{21}x-y}{K_{2}}\right)\end{aligned}$ Linear increase of partner benefit; modification of intrinsic growth. Gause and Witt (1935)
Symbiont–host linear model $\begin{aligned} f(x,y)&=r_{1}\frac{K_{1}+\alpha_{12}y-x}{K_{1}+\alpha_{12}y}\\ g(x,y)&=r_{2}\frac{K_{2}+\alpha_{21}x-y}{K_{2}}\end{aligned}$ Asymmetric symbiont–host structure; linear modification of carrying capacity; obligate for $K_{1}=0$. Whittaker (1970)
Lotka–Volterra mutualism $\begin{aligned} f(x,y)&=r_{1}+\alpha_{12}y-\alpha_{11}x\\ g(x,y)&=r_{2}+\alpha_{21}x-\alpha_{22}y\end{aligned}$ Classical linear mutualistic enhancement of per-capita growth; obligate when $r_{1}/\alpha_{11}\leq 0$. Vandermeer and Boucher (1978)
Multiplicative linear benefit $f(x,y)$ $=r_{1}\left(1-\dfrac{x}{K_{1}}\right)\left(1+\dfrac{\alpha_{12}y}{K_{1}}\right)$ $g(x,y)$ $=r_{2}\left(1-\dfrac{y}{K_{2}}\right)\left(1+\dfrac{\alpha_{21}x}{K_{2}}\right)$ Partner increases effective carrying capacity multiplicatively; facultative. Addicott (1981)
Table 1: Linear–benefit mutualism models expressed through per-capita growth functions $f$ et $g$.
2.1.2 Saturating Mutualism Models (Interspecific Saturation)
Table 2 compiles mutualism models where the benefits gained by a specie is a saturating function of the density of the other species.
Model Per-capita growth functions Interpretation Reference
Birth limitation reduced by partner $\begin{aligned} f(x,y)&=r_{1}-\frac{b_{1}x}{1+m_{1}y}-d_{1}x\\ g(x,y)&=r_{2}-\frac{b_{2}y}{1+m_{2}x}-d_{2}y\end{aligned}$ Partner reduces density-dependent birth limitation; saturating effect with partner density; facultative. Wolin and Lawlor (1984)
Saturating forager mutualism $\begin{aligned} f(x,y)&=r_{1}-c_{1}x+\frac{b_{12}a_{12}y}{1+a_{12}h_{12}y}\\ g(x,y)&=r_{2}-c_{2}y+\frac{b_{21}a_{21}x}{1+a_{21}h_{21}x}\end{aligned}$ Handling time limits mutualistic gains; stabilizes coexistence; fits pollination/foraging systems. Wright (1989)
Exponential saturation of benefits $f(x,y)$ $=r_{10}+(r_{11}-r_{10})$ $\times\left(1-e^{-k_{1}y}\right)-a_{1}x$ $g(x,y)$ $=r_{20}+(r_{21}-r_{20})$ $\times\left(1-e^{-k_{2}x}\right)-a_{2}y$ Benefits saturate exponentially with partner density. Graves et al. (2006)
Plant–pollinator with shared saturation $\begin{aligned} f(x,y)&=\frac{\eta\alpha y}{1+\alpha x+\alpha\beta y}-b-cx\\ g(x,y)&=\frac{\mu\alpha x}{1+\alpha x+\alpha\beta y}-d\end{aligned}$ Shared saturation of benefits; obligate mutualism. Fishman and Hadany (2010)
Table 2: Mutualism models with interspecific saturation of benefits expressed through per-capita growth functions $f$.
2.1.3 Non-Monotonic Mutualism Models and Consumer-Resource Approach
Table 3 compiles mutualism models where the benefits gained by a specie is a non-monotonic function of the density of the other species.
Model Per-capita growth functions Interpretation Reference
Quadratic interaction $\begin{aligned} f(x,y)&=r_{1}\left(c_{1}-x-a_{1}(y-b_{1})^{2}\right)\\ g(x,y)&=r_{2}\left(c_{2}-y-a_{2}(x-b_{2})^{2}\right)\end{aligned}$ Benefit maximized at intermediate partner density; allows transitions between mutualism and competition. Zhang (2003)
Plant–mycorrhizae with cost $\begin{aligned} f(x,y)&=r_{1}\frac{K_{1}+\alpha_{12}y-x}{K_{1}+\alpha_{12}y}-ay\\ g(x,y)&=r_{2}\frac{K_{2}+\alpha_{21}x-y}{K_{2}}\end{aligned}$ Interspecific asymmetric cost; facultative mutualism. Neuhauser and Fargione (2004)
Protection mutualism with saturating mortality reduction $f(x,y)$ $=\rho_{1}b_{1}\left(1-\frac{x}{S_{1}}\right)$ $-\left(d_{1\min}+\frac{d_{1\max}-d_{1\min}}{1+c_{1}y}\right)$ $g(x,y)$ $=\rho_{2}b_{2}\left(1-\frac{y}{S_{2}+x}\right)$ $-\left(d_{2\min}+\frac{d_{2\max}-d_{2\min}}{1+c_{2}x}\right)$ Protection reduces mortality at low partner density; net effects may shift with density. Thompson et al. (2006)
Consumer–resource mutualism $\begin{aligned} f(x,y)&=r_{1}+c_{1}\frac{a_{12}y}{h_{2}+y}\\ &-q_{1}\frac{\beta_{12}y}{e_{1}+x}-s_{1}x\\ g(x,y)&=r_{2}+c_{2}\frac{a_{21}x}{h_{1}+x}\\ &-q_{2}\frac{\beta_{21}x}{e_{2}+y}-s_{2}y\end{aligned}$ Bidirectional consumer–resource structure; obligate when $r_{1}=0$. Holland and DeAngelis (2010)
Ant–fungal garden mutualism $\begin{aligned} f(x,y)&=r_{f}\frac{ay^{2}}{b+ay^{2}}-r_{c}y-d_{1}x\\ g(x,y)&=r_{a}x-d_{2}y\end{aligned}$ Consumer–resource mutualism with nonlinear resource production. Kang et al. (2011)
Plant–mycorrhizae with variable costs $f(x,y)$ $=r_{p}+q_{hp}\frac{\alpha y}{d+x}$ $-q_{cp}\beta y-\mu_{p}x$ $g(x,y)$ $=q_{cm}\beta x-q_{hm}\frac{\alpha x}{d+x}-\mu_{m}y$ Explicit separation of benefit and cost terms; obligate mutualism. Martignoni et al. (2020)
Plant–pollinator consumer–resource model $f(x,y)$ $=b_{P}\left(f+\phi\frac{axy}{1+ahx+axy}\right)$ $-s_{P}x-d_{P}$ $g(x,y)$ $=b_{A}+\epsilon\frac{ax}{1+ahx}-s_{A}y-d_{A}$ Mechanistic pollination model. Hale and Valdovinos (2021)
Table 3: Non-monotonic mutualism and consumer-resource models with peak benefit at intermediate partner density expressed through per-capita growth functions $f$.
@@ 2.2 Key Assumptions and Definition
We have so far not explicitly defined what makes a mutualistic model different than a model considering predator–prey or competitive interactions. The system of differential equations (1) does not, by itself, specify the nature of the interactions, as it also includes classical models such as the Lotka–Volterra predator–prey system. We introduce hereafter definitions of mutualism, parasitism, competition, each imposing specific constraints on the functional forms of $f$ and $g$.
To analyze and interpret the behavior of the general system (1) in a mutualistic context, we follow an approach similar to Brauer and Castillo-Chavez (2012), relying on the partial derivatives of the growth functions. These derivatives capture how the growth rate of each species responds to changes in the density of both species. For this reason, starting from Section 3.4, we will assume that $f$ and $g$ are $\operatorname{\mathcal{C}^{1}}$ functions i.e. continuous and continuously differentiable. We focus on specific regions of the phase plane defined by the population densities $x$ and $y$. In each region, the signs of the partial derivatives of the interaction terms determine the local effect of one species on the growth rate of the other. As a result, an interaction may be positive in one part of the phase portrait and negative in another, depending on the local population densities. Different regions of the phase portrait may correspond to different interaction regimes, described below:
@@
  • Intraspecific Cooperation: $\frac{\partial f}{\partial x}>0$ and $\frac{\partial g}{\partial y}>0$ describe positive feedback within each species, representing cooperative interactions, where individuals of the same population support each other’s growth. This mechanism is also called ’Allee effect’ when the positive feedback occurs at low density.
@@
In contrast to many pre-existing models where these partial derivatives always remain within the same strict sign or even remain constant regardless of species densities (Hale and Valdovinos, 2021), we allow the sign of these partial derivatives to change. Such assumption allows for transitions from mutualistic (positive value) to parasitic interactions (negative value) between species, depending on their respective densities. We thus give an extended definition of mutualism as follows:
@@
These definitions are illustrated in Figure 1, which presents two different situations, both consistent with mutualism as defined in Definition 2.1. In the first case, a strictly mutualistic region appears, although the condition does not hold for all densities $x$ and $y$. In the second case, the system remains mutualistic overall, but no strictly mutualistic region exists; that is, both populations may benefit from the interaction, though not simultaneously.
@@
Figure 1: Illustration of density-dependent effects of mutualism (Definition 2.1) through the signs of cross partial derivatives. In both situations, the interaction is classified as mutualistic according to Definition 2.1, even though the local effects of one species on the other vary. Figure 1(a): $\frac{\partial f}{\partial y}>0$ everywhere, meaning that species $y$ always has a positive effect on the growth of species $x$. The hatched region indicates where $\frac{\partial g}{\partial x}>0$, so that species $x$ also benefits species $y$ locally. This region represents a strictly mutualistic zone. Outside this zone, the effect of species $x$ on $y$ is negative, yet the system still qualifies as mutualistic under Definition 2.1. Figure 1(b): The cross partial derivatives $\frac{\partial f}{\partial y}$ and $\frac{\partial g}{\partial x}$ vary and are positive in distinct, non-overlapping regions: the hatched (resp. dashed) region corresponds to $\frac{\partial g}{\partial x}>0$ (resp. $\frac{\partial f}{\partial y}>0$). Outside these regions, the interaction is parasitic for both species.